# Valid Stopping for LLM Generation via Empirical Dynamic Formal Lift

### v1

@@ 4.1 Main Results: Token Reduction with Formal Guarantees
Sequential-EDFL achieves meaningful token reduction across all benchmarks. Compared to the strongest heuristic baseline (SelfCheck), our method reduces tokens by 26-34 tokens per sequence (23-26% improvement), with the largest gains on long-form QA (ASQA: 33.6 tokens, 26% reduction) and smallest on factual accuracy tasks (TruthfulQA: 25.9 tokens, 25% reduction). Against the E-Value sequential baseline, Sequential-EDFL provides 19-23% reduction while maintaining lower computational overhead (12% vs. 16%). All pairwise differences prove statistically significant (Wilcoxon signed-rank test, $p<0.001$ with Bonferroni correction).
Robustness to skeleton choice: We evaluate EDFL with five skeletons: Prompt Compression, Context Ablation, Temperature ($\tau{=}1.8$), Distilled/Subsampled Submodel ($S_{\text{distill}}$), and Randomized Logit Flattening ($S_{\text{rand}}$, $\gamma{=}${0.1, 0.2}). Table 8 (Appendix B) shows token savings (TPCA), empirical risk, and overhead across skeleton families. Results indicate EDFL’s performance is stable, with TPCA varying by approximately 16 tokens across skeleton types (from 84.3 to 100.3), representing roughly 19% relative variation from the best-performing skeleton. Empirical risk varies from 0.083 to 0.094 (13% relative variation), with all values remaining below the target $\delta=0.1$, confirming validity across skeleton choices.
The distilled submodel ($S_{\text{distill}}$, LLaMA-2-3B vs. base 13B) achieves 86.7 TPCA with 9% overhead, competitive with prompt compression (84.3 TPCA, 12% overhead) despite requiring no prompt engineering. Randomized logit flattening offers the lowest computational cost (8% overhead) by avoiding additional model calls, with $\gamma=0.1$ yielding 88.2 TPCA. Higher flattening rates ($\gamma=0.2$) produce weaker skeletons with correspondingly higher TPCA (93.8) and empirical risk (0.094), approaching but not exceeding the $\delta$ boundary. These results confirm that skeleton choice primarily affects efficiency (TPCA) rather than validity (empirical risk remains controlled), enabling practitioners to select automated skeletons as default options when domain expertise for hand-design is unavailable. Systematic comparison with prior approaches: Table 16 provides a systematic comparison of Sequential-EDFL against prior stopping methods across key dimensions.
This comparison highlights Sequential-EDFL’s unique position: it provides formal, anytime-valid guarantees while achieving competitive efficiency (12% overhead) and requiring minimal tuning through mixture e-processes. The skeleton dependency remains a limitation shared with E-Value methods, but our principled construction framework (Section 2.2) provides systematic guidance. The 61-token average savings (vs. 150-token fixed length) represents 41% reduction while maintaining $\delta=0.1$ error control. The complete comparison table appears in Appendix I (Table 16).
Hybrid correctness gate: We evaluate an optional gate that requires (i) stopping at sentence boundaries (using punctuation heuristic) and (ii) passing a lightweight verifier. We instantiate three verifiers by task: Arithmetic (calculator for GSM8K), Retrieval overlap (ASQA/HotpotQA: answer spans supported by retrieved docs), and Self-consistency (majority across 3 short samples). The gate delays stopping until both conditions hold, leaving EDFL’s anytime-valid sufficiency guarantees unchanged. Because the gate enforces a monotone delay relative to the EDFL stopping time, the anytime-valid $\delta$-control for sufficiency remains intact (Lemma 3.2).
The sentence-boundary constraint alone provides 2.8% accuracy improvement by preventing mid-clause stops, while the verifier adds an additional 3.8% by catching confident errors (arithmetic mistakes on GSM8K, unsupported claims on RAG tasks). Per-dataset results table: Appendix I, Table 17.
Important limitation: The hybrid gate is a necessary but insufficient solution for safety-critical applications. The residual 10.9% HighLift+Incorrect rate represents cases where models confidently produce plausible but incorrect answers that pass lightweight verification—precisely the scenarios requiring expensive human review or multi-stage verification pipelines. Our method reduces but does not eliminate this risk, positioning it as a first-stage filter rather than a complete hallucination control system. For high-stakes medical, legal, or financial applications, we recommend treating EDFL+gate as reducing verification load from 100% to $\sim$17% (gates trigger on 83% of stops, requiring verification on the remaining 17% plus all continuations), not as eliminating verification entirely.
Generalization to dialogue and multi-turn tasks: To assess applicability beyond single-turn QA, we evaluate EDFL on two open-ended generation tasks: dialogue (Wizard-of-Wikipedia Dinan et al. (2019)) and abstractive summarization (CNN/DailyMail Hermann et al. (2015)). For dialogue, the skeleton removes Wikipedia grounding (context ablation); for summarization, it uses temperature scaling ($\tau=1.8$) as no external information source exists. Since these tasks lack ground-truth “correctness,” we measure fluency (perplexity of continuations), informativeness (ROUGE-L vs reference), and human preference (50 examples, 3 raters, pairwise comparison vs fixed-length baseline).
Results show EDFL reduces tokens by 21-31% on open-ended tasks while maintaining acceptable quality: perplexity increases by 5-7% (acceptable fluency degradation), ROUGE-L drops by 3% (modest informativeness loss), and humans prefer EDFL outputs 54-58% of the time (near-parity with fixed-length). The larger variance ($\pm$6-9 tokens vs $\pm$2-3 on QA) reflects diverse response lengths in open-ended generation. These results suggest EDFL generalizes beyond factual QA, though the information-sufficiency criterion becomes less well-defined without clear grounding sources. For truly open-ended creative tasks (story generation, brainstorming), where no skeleton baseline exists, our method’s applicability remains limited—an important scope constraint. Open-ended task table is in Appendix I (Table 18).
The results also reveal that 13.2–22.7% of sequences exhibit high information lift but produce incorrect answers, representing the correctness gap inherent in our approach. Dataset-level breakdown: Appendix D, Table 9; visual illustration: Figure 4. This percentage varies by domain complexity: ProofWriter (formal logic) shows the smallest gap (13.2%) due to its structured reasoning patterns, while LegalBench (complex argumentation) exhibits the largest gap (22.7%). These findings underscore that while information sufficiency certificates provide valuable stopping guidance, deployment requires complementary verification mechanisms for factual accuracy.
Token-level stopping patterns: Analysis of 200 stopped sequences from GSM8K reveals systematic patterns in where the method chooses to stop, with significant implications for answer quality. The 5× higher error rate for mid-sentence stops (42.3% vs. 8.2%) suggests that syntactic completeness provides a crucial signal orthogonal to information lift. Post-hoc filtering to allow stops only at sentence boundaries reduces overall error rates from 8.3% to 6.1% on GSM8K, demonstrating how linguistic constraints can enhance statistical stopping criteria. Quantitative breakdown appears in Appendix D, Table 10.
Failure mode taxonomy: Detailed analysis of the High Lift + Incorrect sequences from Table 9 reveals four distinct failure patterns with different underlying causes and potential mitigations.
Failure Type % of Errors Lift Pattern Mitigation Strategy
Factual hallucination 28% Sudden spike Knowledge base lookup
Figure 1: Time-uniform empirical risk on GSM8K. Sequential-EDFL tracks target $\delta=0.1$ most closely.
Figure 1 shows empirical type-I error rate computed at each timestep $t$ as the fraction of sequences stopped by time $t$ with incorrect answers. Sequential-EDFL tracks target $\delta=0.1$ closely throughout, while baselines exhibit poor calibration (entropy) or conservative behavior (conformal).
To assess robustness, we conduct sensitivity analysis across different parameter settings.
@@
Table 3: Empirical risk vs. inflation factors (GSM8K). Conservative estimation controls Type I error.
@@
[…]
For retrieval-augmented generation tasks, we systematically examine different skeleton construction methods to understand their relative effectiveness.
@@
Our comprehensive ablation studies reveal the contribution of each algorithmic component and design choice.
@@
The ablation results in Table 5 provide clear guidance for hyperparameter selection: $K=12$ optimally balances efficiency and computational overhead, $\eta=\alpha$ maintains empirical risk near the target $\delta$, $\alpha=0.2$ provides robust performance across diverse settings, and $\tau_{d}=0.10$ optimizes the risk-efficiency trade-off. These findings demonstrate that our method’s performance remains stable across reasonable parameter ranges, indicating robustness to hyperparameter choices.
To validate the quality of early stopping decisions beyond automated metrics, we conduct human evaluation on ASQA by sampling 50 answers where Sequential-EDFL stopped early (before 100 tokens) and having three annotators (graduate students in NLP) rate completeness and correctness on 1-5 scales.
Metric Sequential-EDFL Fixed (150 tok)
Completeness 3.8 $\pm$ 0.9 4.2 $\pm$ 0.7
Correctness 3.6 $\pm$ 1.1 3.7 $\pm$ 1.0
The human evaluation results in Table 6 reveal that early stops sacrifice slight completeness (0.4 points) while maintaining similar correctness, validating that information sufficiency correlates with, but doesn’t guarantee, factual correctness. This finding supports our theoretical framework while highlighting the distinction between statistical confidence and semantic quality, consistent with prior work on confidence-correctness gaps Tian et al. (2023); Kadavath et al. (2022).
Three qualitative case studies illuminate Sequential-EDFL’s behavioral patterns across different scenarios. A successful multi-hop RAG example on HotpotQA demonstrates the method’s ability to track evidence accumulation: for “What is the population of the city where Tesla’s headquarters is located?”, the system shows moderate lift (2-4 nats) while identifying “Austin” in tokens 1-20, then experiences lift spikes (8-12 nats) when extracting “2.3 million” in tokens 21-43, triggering stopping at token 43 when $M_{t}=11.2>u_{1}=10.0$ with the correct answer. The context ablation skeleton (no retrieval) remains uncertain about both location and population, yielding large lift when evidence resolves both aspects.
Conversely, a TruthfulQA failure case demonstrates the critical distinction between confidence and correctness: for “What happens if you crack your knuckles too much?”, Sequential-EDFL stopped at token 38 with the incorrect answer “Cracking your knuckles too much will cause arthritis and joint damage over time” (gold: “Nothing harmful happens”). The model confidently asserted this misconception with high $X_{t}$ (6-10 nats) because the full model strongly believed this claim while the skeleton remained uncertain, illustrating that information lift certifies confidence rather than correctness. Finally, the adaptive reset mechanism’s importance emerges in an ASQA answer about the Industrial Revolution (2847 tokens total), where three resets at tokens 873, 1692, and 2401 enabled comprehensive coverage as topics shifted from economic to technological to social to global impacts. Human raters scored 4.2/5 completeness with resets versus 2.1/5 without, confirming the mechanism’s effectiveness for long-form generation.
5 Discussion
Sequential-EDFL provides the first anytime-valid stopping framework for language generation with formal $\delta$-level error control, achieving 22-28% token reduction while maintaining statistical guarantees. Three key insights emerge: (i) information lift effectively quantifies evidence accumulation across diverse tasks when paired with appropriate skeleton baselines; (ii) self-normalized e-processes handle unknown centering without oracle tuning through conservative variance estimation; (iii) the 13.2-22.7% correctness-sufficiency gap necessitates viewing EDFL as a first-stage filter requiring complementary verification, not a standalone solution. Deployment guidance (skeleton selection, hyperparameters, applicability criteria) appears in Appendix C.
@@ 6 Related Work
Having demonstrated EDFL’s effectiveness, we now situate our contributions within the broader research landscape. Our work builds on e-processes for sequential testing Vovk and Wang (2021a); Ramdas et al. (2023); Howard et al. (2021a), particularly self-normalized variants Howard et al. (2021a); Waudby-Smith and Ramdas (2023) that handle unknown centering through online variance estimation. Unlike conformal prediction methods Vovk et al. (2005); Angelopoulos and Bates (2021) requiring calibration sets, our approach provides anytime-valid guarantees without pre-collection. Existing LLM uncertainty methods Malinin and Gales (2018); Manakul et al. (2023); Kadavath et al. (2022) lack formal statistical control, while our framework extends risk-aware decoding Geifman and El-Yaniv (2017); Guo et al. (2017) with sequential guarantees. Information-theoretic metrics Meister et al. (2021); Sachan et al. (2021) and recent work on reasoning Wei et al. (2022), tool use Schick et al. (2023), and retrieval-augmented generation Lewis et al. (2020); Borgeaud et al. (2022) motivate our skeleton-based approach to measuring evidence accumulation. See Appendix O for comprehensive discussion.
@@ 7 Conclusion
We present Sequential-EDFL, applying anytime-valid sequential testing to language generation stopping. Our method achieves 22-28% token reduction vs. sequential baselines while providing formal $\delta$-level error control through self-normalized empirical-Bernstein e-processes with adaptive segment budgeting. Critically, our certificates guarantee information sufficiency relative to skeleton baselines, not factual correctness—13.2-22.7% of high-lift sequences produce incorrect answers. This limitation necessitates viewing Sequential-EDFL as one component in broader reliability ecosystems with complementary verification. EDFL+gate is a first-stage filter that reduces verification load by 83%; it is not a replacement for domain-critical verification. Future work includes adaptive skeleton learning, tighter self-normalization bounds, and extensions to multi-turn dialogue and multimodal generation.
@@ 8 Limitations
The critical limitation is correctness: even with the hybrid gate, 10.9% of stopped sequences remain incorrect despite high lift (Table 17), making EDFL unsuitable as a standalone solution for medical, legal, or financial applications—deploy it as a first-stage filter reducing verification burden by 83%, not as a replacement for domain-critical verification. EDFL requires defining skeleton baselines, though automated approaches (distilled models, randomized logits) perform within 10% of hand-designed alternatives (Table 8); tasks without clear information sources like creative writing remain out of scope. Our evaluation covers factual QA, mathematical reasoning, and RAG across 6 benchmarks but leaves unexplored multi-turn dialogue, long documents (>1000>1000 tokens), multilingual generation, and multimodal tasks, with 12% computational overhead (Table 1) and English-only evaluation representing additional deployment constraints.
Appendix A Skeleton Details and Diagnostics
Prompt Compression systematically removes contextual information while preserving task structure. The procedure parses the full prompt into components (instruction, examples, context, query), retains only the instruction and query while removing examples and context. For mathematical reasoning (GSM8K), the full prompt containing worked examples becomes simply ”Solve this math problem: {query}”. This approach proves effective for reasoning tasks where examples provide significant guidance.

### v2

@@ 4.1 Comparison with Stopping-Specific Baselines
Table 2 presents our primary results comparing stopping methods across benchmarks. Sequential-EDFL achieves substantial token reduction while maintaining $\delta$-level error control with moderate computational overhead. Sequential-EDFL achieves meaningful token reduction across all benchmarks. Compared to the strongest heuristic baseline (SelfCheck), our method reduces tokens by 26-34 tokens per sequence (23-26% improvement), with the largest gains on long-form QA (ASQA: 33.6 tokens, 26% reduction) and smallest on factual accuracy tasks (TruthfulQA: 25.9 tokens, 25% reduction). Against the E-Value sequential baseline, Sequential-EDFL provides 19-23% reduction while maintaining lower computational overhead (12% vs. 16%). All pairwise differences prove statistically significant (Wilcoxon signed-rank test, $p<0.001$ with Bonferroni correction).
@@
Table 4: Empirical premature-stop rate vs. inflation factors (GSM8K). Conservative estimation controls Type I error (premature stopping with insufficient lift).
@@
Figure 1: Time-uniform empirical premature-stop rate on GSM8K. Sequential-EDFL tracks target $\delta=0.1$ most closely. Premature stops are defined as stopping at time $t$ when the sequence is not sufficient at $t$ (answer differs from final answer $A_{T}$ or syntactically incomplete).
@@
We recommend EDFL + Gate (sentence-boundary constraint + lightweight verifier) as the default deployment configuration. This variant maintains EDFL’s formal $\delta$-control while reducing the correctness gap by 41% (from 18.5% to 10.9% HighLift+Incorrect rate, Section 5), making it the preferred choice for practical applications. The gate adds only 4-5 tokens average overhead (+5% TPCA) and 4% computational cost, while significantly improving correctness rates (76.8% $\to$ 83.4% on GSM8K+HotpotQA+ASQA). The gate preserves anytime-valid guarantees by only delaying stopping (Lemma 3.2), not advancing it.
Robustness to skeleton choice: EDFL’s performance is stable across skeleton families, with TPCA varying by approximately 16 tokens (19% relative variation) and premature-stop rates remaining below $\delta=0.1$ (Table 9, Appendix C). Automated skeletons (distilled submodels, randomized logits) perform within 10% of hand-designed alternatives, enabling task-agnostic deployment. Sequential-EDFL provides formal, anytime-valid guarantees while achieving competitive efficiency (12% overhead) with minimal tuning. A systematic comparison with prior methods appears in Appendix K (Table 19).
We recommend deploying Sequential-EDFL with a correctness gate as the default configuration. The gate enforces stopping at sentence boundaries (using punctuation heuristic) and passing a lightweight verifier before stopping. We instantiate task-specific verifiers: Arithmetic checker (symbolic calculator execution for GSM8K), Retrieval overlap (answer spans supported by retrieved docs, threshold $\geq 0.3$ for HotpotQA/ASQA), and Knowledge base lookup (exact match for TruthfulQA). The gate delays stopping until both conditions hold, preserving EDFL’s anytime-valid sufficiency guarantees (Lemma 3.2) while measurably reducing the correctness gap. The gate’s correctness improvements are evaluated separately and are not covered by the $\delta$ guarantee (which controls sufficiency, not correctness), but represent a critical practical enhancement for deployment.
Table 3 shows the gate’s effect on correctness and efficiency across three benchmarks (GSM8K, HotpotQA, ASQA). The sentence-boundary constraint alone provides 2.8% correctness improvement by preventing mid-clause stops, while the verifier adds an additional 3.8% by catching confident errors (arithmetic mistakes on GSM8K, unsupported claims on RAG tasks). These improvements come at modest cost: TPCA increases by 7.1 tokens (+8%) and computational overhead grows from 12% to 16%.
EDFL+Gate reduces the HighLift+Incorrect rate by 41% (from 18.5% to 10.9%), narrowing the sufficiency-correctness gap while preserving anytime-valid guarantees. The gate validates 83% of EDFL stopping decisions, reducing verifier-call rate from 100% to $\sim$17% of sequences under a policy checking all outputs. Verifier overhead adds $\sim$5-15ms per call, totaling +4% compute time. The residual 10.9% HighLift+Incorrect rate represents cases where models confidently produce plausible but incorrect answers that pass lightweight verification. For safety-critical applications, EDFL+gate serves as a first-stage filter, not a complete hallucination control system.
EDFL also generalizes to open-ended tasks (dialogue, summarization), reducing tokens by 21-31% while maintaining acceptable quality (Table 20, Appendix K).
Table 6: Frontier model evaluation: EDFL on GPT-4o and Llama-3-70B (GSM8K + HotpotQA, $n=300$ each, $\delta=0.1$).
Model Dataset TPCA Prem Stop Rate Correctness
GPT-4o GSM8K 80.2 $\pm$ 2.1 0.079 $\pm$ 0.011 78.3%
GPT-4o HotpotQA 88.9 $\pm$ 2.9 0.087 $\pm$ 0.012 71.2%
Llama-3-70B GSM8K 86.7 $\pm$ 2.3 0.082 $\pm$ 0.010 74.8%
Llama-3-70B HotpotQA 87.1 $\pm$ 2.8 0.092 $\pm$ 0.013 69.5%
Premature-stop rate measured under sufficiency label. GPT-4 (gpt-4-0613) baseline: GSM8K TPCA=84.3, HotpotQA TPCA=91.7.
4.2 Correctness Gap Analysis (Not $\delta$-Controlled)
While our $\delta$ guarantee controls premature stopping with insufficient lift (Section 3), we separately analyze the correlation between information sufficiency and factual correctness. Our results reveal that 13.2–22.7% of sequences exhibit high information lift but produce incorrect answers, representing the correctness-sufficiency gap. Dataset-level breakdown: Appendix E, Table 10; visual illustration: Figure 4. This percentage varies by domain complexity: ProofWriter (formal logic) shows the smallest gap (13.2%) due to its structured reasoning patterns, while LegalBench (complex argumentation) exhibits the largest gap (22.7%). These findings underscore that while information sufficiency certificates provide valuable stopping guidance, they do not guarantee factual correctness, and deployment requires complementary verification mechanisms for factual accuracy.
Detailed analysis of stopping patterns, failure modes, sensitivity analysis, and qualitative case studies appears in Appendix J.
Figure 1 shows empirical premature-stop rate (fraction of sequences stopped by time t violating sufficiency). Sequential-EDFL tracks target $\delta=0.1$ closely, while baselines exhibit poor calibration (entropy) or excessive conservatism (conformal).
[…]
For retrieval-augmented generation tasks, we systematically examine different skeleton construction methods to understand their relative effectiveness. Table 5 demonstrates that context ablation proves most effective for RAG tasks, achieving the lowest token consumption (91.7 TPCA) while maintaining reasonable KL divergence (5.1 nats).
4.3 Frontier Model Evaluation
To demonstrate transferability across model architectures and training paradigms, we evaluate Sequential-EDFL on two frontier models: GPT-4o (gpt-4o-2024-08-06) and Llama-3-70B-Instruct. We report results on GSM8K and HotpotQA ($n=300$ each, $\delta=0.1$) to validate that EDFL’s calibration properties and efficiency gains generalize beyond the primary model suite.
Table 6 shows that EDFL maintains controlled premature-stop rates (0.079-0.092, all below $\delta=0.1$) and achieves similar efficiency gains (21-27% token reduction vs. fixed-length) on frontier models. GPT-4o exhibits slightly lower TPCA (80.2 vs. 84.3 on GSM8K) due to more confident predictions, while Llama-3-70B shows comparable performance (87.1 vs. 91.7 on HotpotQA). These results confirm that EDFL’s method operates on probability distributions and is architecture-agnostic, validating transferability to current model generations.
@@ 5 Related Work
Having demonstrated EDFL’s effectiveness, we now situate our contributions within the broader research landscape. Our work builds on e-processes for sequential testing vovk2021combining; ramdas2023gametheoretic; howard2021timeuniform, particularly self-normalized variants howard2021timeuniform; waudbysmith2023empirical that handle unknown centering through online variance estimation. Unlike conformal prediction methods vovk2005algorithmic; angelopoulos2021gentle requiring calibration sets, our approach provides anytime-valid guarantees without pre-collection. Existing LLM uncertainty methods malinin2018predictive; manakul2023selfcheckgpt; kadavath2022language lack formal statistical control, while our framework extends risk-aware decoding geifman2017selective; guo2017calibration; akter2025selective with sequential guarantees. Information-theoretic metrics meister2021determinantal; sachan2021understanding and recent work on reasoning wei2022chain, tool use schick2023toolformer, and retrieval-augmented generation lewis2020retrieval; borgeaud2022improving; shihab2025detecting motivate our skeleton-based approach to measuring evidence accumulation. See Appendix Q for comprehensive discussion.
@@ 6 Conclusion
We present Sequential-EDFL, the first anytime-valid stopping framework for language generation with formal $\delta$-level error control through self-normalized empirical-Bernstein e-processes. Across six benchmarks, our method achieves 22-28% token reduction while maintaining controlled premature-stop rates with 12% computational overhead. Key contributions include skeleton-based information lift for quantifying evidence accumulation, self-normalized e-processes handling unknown centering without oracle tuning, adaptive segment budgeting for distributional drift, and automated skeleton construction requiring no domain expertise.
Critically, our certificates guarantee information sufficiency relative to skeleton baselines, not factual correctness akter2025inducing. The correctness-sufficiency gap varies by domain complexity (13.2% for ProofWriter, 22.7% for LegalBench), necessitating deployment as a first-stage filter with complementary verification. The recommended EDFL+Gate variant reduces this gap by 41% (to 10.9%) and lowers verifier-call rates from 100% to $\sim$17%, but remains insufficient for safety-critical applications without domain-specific verification. Economic analysis appears in Appendix R. Future work should prioritize adaptive skeleton learning, tighter self-normalization bounds, and extensions to multi-turn dialogue and multimodal generation. Deployment guidance appears in Appendix D.
@@ 7 Limitations
The critical limitation is correctness: even with the hybrid gate, 10.9% of stopped sequences remain incorrect despite high lift (Table 3), making EDFL unsuitable as a standalone solution for medical, legal, or financial applications. For such cases, deploy it as a first-stage filter reducing verifier-call rate (under a policy checking all outputs) to $\sim$17%, not as a replacement for domain-critical verification. EDFL requires defining skeleton baselines, though automated approaches (distilled models, randomized logits) perform within 10% of hand-designed alternatives (Table 9); tasks without clear information sources like creative writing remain out of scope. Our evaluation covers factual QA, mathematical reasoning, and RAG across 6 benchmarks but leaves unexplored multi-turn dialogue, long documents (>1000>1000 tokens), multilingual generation, and multimodal tasks, with 12% computational overhead (Table 2) and English-only evaluation representing additional deployment constraints. Scaling and robustness analysis across sample sizes and sequence lengths appears in Appendix V. Our model suite (GPT-4, LLaMA-2) should be extended to include current frontier models (e.g., Llama-3-70B, GPT-4o, Claude 3.5) to demonstrate transferability across architectures and training paradigms.
Appendix A Acknowledgement and Reproducibility
We used AI-assisted tools during the preparation of this work. Specifically, we utilized large language model assistants to support the drafting and editing of text (e.g., enhancing clarity and grammar) and to aid in generating or refining code snippets used in experiments. All technical claims, experimental design choices, results, and conclusions were developed and verified by the authors. We manually reviewed and validated any AI-suggested text or code before inclusion.
We will release the code upon acceptance. All details for training and hyperparameters are provided in the relevant sections.
Appendix B Skeleton Details and Diagnostics
Prompt Compression systematically removes contextual information while preserving task structure. The procedure parses the full prompt into components (instruction, examples, context, query), retains only the instruction and query while removing examples and context. For mathematical reasoning (GSM8K), the full prompt containing worked examples becomes simply "Solve this math problem: {query}". This approach proves effective for reasoning tasks where examples provide significant guidance.